1. Considering the total number of electrons in NO2, why might it be expected that the reaction of a free radical with NO2 is a chain-terminating reaction?
  2. The average atmospheric pressure at sea level is 1.012 * 106 dynes/m2. The value of g (acceleration of gravity) at sea level is 980 cm / sec2. What is the mass in kg of the column of air having a cross-sectional area of 1.00 cm2 at the earth's surface and extending to the limits of the atmosphere? (Recall that the dyne is a unit of force, and force = mass * acceleration of gravity.
  3. Suppose that 22.4 liters of air at STP is used to burn 1.50g of carbon to form CO2 and that gaseous product is adjusted to STP. What is the volume and average molar mass of the resulting mixture?
1
Expert's answer
2021-10-21T02:09:25-0400

The equation is CO2 + C → 2CO (G)

one mole of CO2 produces 2 moles of CO.

one mole of any gas at STP occupies 22.4L

so two moles of CO will occupy 22.4 × 2 = 44.8L.

so P = 44.8L.

so 2P/10 = 2×44.8/10 = 8.96.